# Thompson sampling for zero-inflated count outcomes with an application to the Drink Less mobile health study

### v1

@@ 7.3 Simulation Environment
We generate $100$ independent simulated datasets using the aforementioned model, evaluating the algorithms based on their cumulative regret. In each simulated experiment, we randomly draw $N=50$ users, with each user participating in the study over $200$ days. We designate an initial exploration period ($\tau=20$) for each algorithm to ensure adequate exploration before exploitation begins.
7.4 Results
@@
Figure 4 illustrates the cumulative regret averaged over $50$ users and $100$ iterations of six different algorithms. This figure reveals that initially, the Linear TS algorithm with a logarithmic transformation of the outcome exhibits lower regret. However, as the experiment progresses and more data is collected, the newly introduced TS-Count variants demonstrate superior performance in minimizing regret compared to the Linear TS. This delayed efficacy may be due to the TS-Count variants requiring more data for learning so that their enhanced modeling capabilities become effective. The uniform strategy, which assigns notifications with a constant and balanced probability, shows a consistently linear increase in regret due to its independence on user feedback. Overall, our findings suggest that the proposed approaches significantly improve user engagement with the app.
@@
Figure 4: Results of the compared algorithms under the setup of the Drink Less study. The results are calculated from $100$ replications of the experiment. The solid lines indicate the mean values, while the shaded bands represent the standard error bounds across the independent replications.
@@ 8 Discussion
In the realm of mHealth, delivering just-in-time adaptive interventions is crucial for improving the distal outcome, which often pertains to the clinical condition of users. To achieve this goal, contextual bandits provide a suitable framework for tailoring mHealth interventions. However, in real-world scenarios, such as the Drink Less study (Bell et al., 2020), proximal outcomes are often collected as counts. This paper addresses the challenge of optimizing intervention delivery in mHealth applications with count-valued proximal outcomes. Specifically, we propose Bayesian sequential approaches that combine common count data models with the Thompson sampling strategy. This combination effectively addresses the exploration and exploitation trade-off in optimizing intervention delivery that targets a count proximal outcome. In the Drink Less study, we find that deploying the proposed TS-Count approaches can significantly improve engagement with the app, which has an important impact in practice since it would benefit current users within the trial.
@@
Within this paper, we have conducted comparisons across a suite of count models—including Poisson, NB, ZIP, and ZINB—considering varying degrees of overdispersion and zero-inflation within an online contextual bandit framework. Interestingly, our findings suggest that more complex models like the NB and ZINB regressions, which offer mechanisms to accommodate overdispersion unlike their Poisson equivalents, may not outperform in an online environment. This could be attributed to the limited data typically available online and the complexity introduced by the additional dispersion parameter that necessitates estimation. Another reason might be that, despite overdispersion, the Poisson model can still yield consistent parameter estimates which are adequate for bounding the regret. Nevertheless, properly addressing zero-inflation is critical, as failure to do so can result in biased estimates and linear regret.
@@
Previous studies on contextual bandits based on generalized linear models have focused on a generic environment for online decision-making (Filippi et al., 2010; Li et al., 2017). However, they left unresolved statistical issues in applying contextual bandits to mHealth studies, including how to model overdispersed and zero-inflated count proximal outcomes in an online setting and whether a proper statistical model can improve the performance of online decision-making. Our work addresses these issues by providing both theoretical guarantees and evaluating their empirical performances. These algorithms, however, are not limited to mHealth studies and can be applied to other studies with a count outcome.
@@
There are opportunities to extend this work in several directions. For instance, many mHealth studies involve multiple users being intervened at each decision time. By incorporating mechanisms for information borrowing between users, as suggested by Tomkins et al. (2021), our algorithms could learn more efficiently and accommodate user heterogeneity. Additionally, exploring alternative approximation methods for sampling from posterior distributions presents a substantial opportunity for improvement. Techniques such as stochastic gradient descent (Ding et al., 2021) and Langevin Monte Carlo methods (Xu et al., 2022) could be adapted for use in zero-inflated models. These methods promise more computationally efficient algorithms, a crucial factor in large-scale and real-time applications. Furthermore, a straightforward extension would be differentiating the feature vectors of the zero component and the Poisson component in ZIP and ZINB models. This modification would allow the models to more accurately reflect real-world scenarios where these components are likely to differ.
@@ A.1 Proof of Bayesian regret for TS-NB
$\leq$ $\tau\Delta_{\max}+\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[\Delta_{A_{t}}\right]$
$=$ $\tau\Delta_{\max}+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[R_{t}\left(A_{t}^{*},\beta^{*}\right)-U_{t}\left(A_{t}^{*},\hat{\beta}_{t}\right)\right]}_{\mathcal{R}_{\text{Bayes }}^{(1)}(T)}+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[U_{t}\left(A_{t},\hat{\beta}_{t}\right)-R_{t}\left(A_{t},\beta^{*}\right)\right]}_{\mathcal{R}_{\text{Bayes}}^{(2)}(T)}.$
@@
where $\mathcal{O}(1)$ comes from the failure event of the concentration of $\hat{\beta}_{t}$ in Lemma A.1. According to Lemma 2 in Li et al. [2017], as Assumption 1 in the main manuscript and $\lambda_{\min}(G_{\tau})\geq 1$ are satisfied, we have
@@
$\mathcal{R}_{\text{Bayes}}(T)\leq\tau\Delta_{\max}+\mathcal{O}(1)+\left[2sY_{\max}\kappa_{\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)}.$
@@ A.2 Proof of Bayesian regret for TS-ZIP
$\leq$ $\tau\Delta_{\max}+\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[\Delta_{A_{t}}\right]$
$=$ $\tau\Delta_{\max}$
$+$ $\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[R_{t}\left(A_{t}^{*},\gamma^{*},\beta^{*}\right)-U_{t}\left(A_{t}^{*},\hat{\gamma}_{t},\hat{\beta}_{t}\right)\right]}_{\mathcal{R}_{\text{Bayes }}^{(1)}(T)}+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[U_{t}\left(A_{t},\hat{\gamma}_{t},\hat{\beta}_{t}\right)-R_{t}\left(A_{t},\gamma^{*},\beta^{*}\right)\right]}_{\mathcal{R}_{\text{Bayes}}^{(2)}(T)}.$
@@
$*\displaystyle(\|\phi(A_{t},X_{t})\hat{\gamma}-c_{1}^{\gamma}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}}-\phi(A_{t},X_{t})\gamma^{*}\|$
@@
$\leq$ $q(2c_{1}^{\gamma}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}}+2c_{1}^{\beta}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}})$
@@
$\mathcal{R}_{\text{Bayes }}(T)$$\leq$ $\tau\Delta_{\max}+\mathcal{O}(1)+\left[2qY_{\max}\kappa_{1,\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)}$
@@ B.1 Proof of frequentist regret for TS-NB
$\mathbb{E}_{t}\left[\Delta_{A_{t}}\right]\leq s\left(c_{1}+c_{2}\right)\left(1+\frac{2}{p_{3}-p_{2}}\right)\times\mathbb{E}_{t}\left[\left\|\phi(A_{t},X_{t})\right\|_{G_{t}^{-1}}\right]+\Delta_{\max}p_{2}.$
@@
where $\alpha$ and $c_{2}$ are set as in Lemma A.4. For these settings, $p_{2}\leq 1/T$ and $p_{3}\geq 0.15$. Finally, to bound $\sum_{t=\tau+1}^{T}\left\|\phi(A_{t},X_{t})\right\|_{G_{t}^{-1}}$, we use Lemma 2 in Li et al. [2017]. As a result, we have
@@ B.2 Proof of frequentist regret for TS-ZIP
$+\Delta_{\max}(\mathbb{P}_{t}(\bar{E}_{2,t}^{\beta})+\mathbb{P}_{t}(\bar{E}_{2,t}^{\gamma})-\mathbb{P}_{t}(\bar{E}_{2,t}^{\beta})\mathbb{P}_{t}(\bar{E}_{2,t}^{\gamma})).$
@@
By setting $\alpha^{\beta}$, $\alpha^{\gamma}$, $c_{2}^{\beta}$, and $c_{2}^{\gamma}$ as in Lemma B.4, we have $p_{2}^{\beta}\leq 1/T$, $p_{2}^{\gamma}\leq 1/T$, $p_{3}^{\beta}\geq 0.15$, and $p_{3}^{\gamma}\geq 0.15$. Finally, to bound $\sum_{\tau+1}^{T}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}}$, we use Lemma 2 in Li et al. [2017]. As a result, we have
@@ Appendix C Technical Lemmas
Using the definition of $S_{t}(\beta^{*})$ and the fact that $\frac{r}{r+\exp(X_{i,A_{i}}^{\top}\beta^{*})}<1$, we can show that $\|S_{t}(\beta^{*})\|_{G_{t}^{-1}}\leq\|Z_{t}\|_{G_{t}^{-1}}$. To bound $\|Z_{t}\|_{G_{t}^{-1}}$, we can use Theorem 1 in Abbasi-Yadkori et al. [2011], which states that if the noise is sub-Gaussian with parameter $\sigma$ (in our case, $\sigma=Y_{\max}$), then with probability at least $1-\delta$, we have
@@
By Lemma 11 in Abbasi-Yadkori et al. [2011] and Assumption 1 from the main manuscript, we have $\log\det(G_{t})\leq d\log(T/d)$. By the choice of $\tau$, $\det(G_{\tau})^{-1}\leq 1$. It follows that

### v2

@@ 7.3.1 Simulation Environment
We generate $100$ independent simulated datasets using the aforementioned models, evaluating the algorithms based on their cumulative regret. In each simulated experiment, we randomly draw $N=349$ users, with each user participating in the study for over $200$ days. Additional simulations on scenarios with $p_{\min}=0$ and $p_{\max}=1$ and user-specific performances are included in Section S4.1 of the Supplement.
7.3.2 Results
@@
Figure 5 illustrates the cumulative regret averaged over $349$ users and $100$ iterations of six different algorithms. The static strategy exhibits a consistently linear increase in regret, attributed to its lack of responsiveness to user feedback. As shown in Figure 5a, the Linear TS algorithm with a logarithmic transformation initially demonstrates lower regret. However, with the accumulation of data throughout the experiment, the TS-Count variants begin to outperform the Linear TS, indicating superior performance in minimizing regret. This delayed efficacy may be due to the TS-Count variants requiring more data for learning so that their enhanced modeling capabilities become effective. Figure 5b further highlights that in the early stages of the trial (where $T\leq 40$), Linear TS performs better than TS-ZIP, aligning with findings from Trella et al. (2022) comparing these algorithms.
Overall, the findings suggest that the proposed approaches, particularly TS-Poisson, significantly improve user engagement with the app, particularly when $T$ is relatively large. For scenarios with a small $T$, Linear TS with a logarithmic transformation could still serve as a good learning algorithm.
(a) OP
(b) ZIOP
@@
Figure 5: Results of the compared algorithms under the Drink Less study with $p_{\min}=0.01$ and $p_{\max}=0.99$. The results are calculated from $100$ replications of the experiment. The solid lines indicate the mean values, while the shaded bands represent the standard error bounds across the independent replications.
7.4 Off-Policy Evaluation on Drink Less
Next, we use off-policy evaluation methods to evaluate the candidate algorithms directly using real data from the Drink Less MRT. We demonstrate that the proposed algorithms lead to greater user-averaged expected rewards.
7.4.1 Methods
We applied the proposed algorithms to the real Drink Less MRT data, utilizing the method of Li et al. (2010). Specifically, at each time $t$, given the collected history $D_{t}$, if the intervention selected by the algorithm matches the intervention assigned in Drink Less, the data tuple $(X_{t},A_{t},Y_{t})$ is retained, and the algorithm is updated. Conversely, if the intervention selected by the algorithm differs from the intervention assigned in Drink Less, the data tuple is disregarded, and the algorithm proceeds to the next decision time.
Using the collected data, we employ the self-normalized inverse probability weighting method to form an unbiased estimate of the per-trial expected reward achieved by each algorithm (Huch et al., 2023):
$\widehat{R}=\sum_{t=1}^{T}\frac{\frac{\pi_{t}(a_{t}|X_{t})}{p_{t}(a_{t}|X_{t})}Y_{t}}{\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|X_{t})}{p_{t}(a_{t}|X_{t})}},$
where $\pi_{t}(a_{t}|X_{t})=0.6$ represents the actual randomization probability in Drink Less and $p_{t}(a_{t}|X_{t})$ represents the probability generated by the algorithm.
We compute the reward improvement, defined as the expected reward of the algorithm minus the expected reward under the pre-specified Drink Less randomization strategy, for each candidate algorithm. We perform this evaluation using $200$ bootstrap samples, running the algorithms separately for each user in each sample. The final results are averaged across all users and datasets.
7.4.2 Results
We present the per-trial expected reward improvement of five algorithms in Figure 6, which demonstrates that TS-Poisson outperforms the other algorithms, with TS-NB as the next most effective. However, the brief duration of the Drink Less MRT data ($T=30$) may compromise the reliability of these expected reward estimates.
Figure 6: Boxplots of the estimates of the reward improvement for five competing algorithms. The results are averaged over $200$ Bootstrap samples.
@@ 8 Discussion
Contextual bandits provide a suitable framework for tailoring mHealth interventions. However, in real-world scenarios, such as the Drink Less study (Bell et al., 2020), proximal outcomes are often collected as counts. This paper addresses the challenge of optimizing intervention delivery in mHealth applications with count proximal outcomes. Specifically, we propose methods that combine common count data models with the Thompson sampling strategy. We provide theoretical analyses on both standard regret and clipped regret for the proposed algorithms.
@@
Within this paper, we have conducted comparisons across a suite of count models—including Poisson, NB, ZIP, and ZINB—considering varying degrees of overdispersion and zero-inflation within an online contextual bandit framework. Interestingly, our findings suggest that more complex models like the NB and ZINB regressions, which offer mechanisms to accommodate overdispersion unlike their Poisson counterparts, may not outperform in an online environment. This could be attributed to the limited data typically available online and the complexity introduced by the additional dispersion parameter that necessitates estimation. Another reason might be that, despite overdispersion, the Poisson model can still yield consistent parameter estimates which are adequate for bounding the regret.
@@
Previous studies on contextual bandits based on generalized linear models have focused on a generic environment for online decision-making (Filippi et al., 2010; Li, Lu and Zhou, 2017). However, they left unresolved statistical issues in applying contextual bandits to mHealth studies, including how to model overdispersed and zero-inflated count proximal outcomes in an online setting and whether a proper statistical model can improve the performance of online decision-making. Our work addresses these issues by providing both theoretical guarantees and evaluating their empirical performances. These algorithms, however, are not limited to mHealth studies and can be applied to other studies with a count outcome.
@@
There are opportunities to extend this work in several directions. For instance, many mHealth studies involve multiple users being intervened at each decision time. By incorporating mechanisms for information borrowing between users, as suggested by Tomkins et al. (2021), our algorithms could learn more efficiently and accommodate user heterogeneity. Additionally, exploring alternative approximation methods for sampling from posterior distributions presents a substantial opportunity for improvement. Techniques such as stochastic gradient descent (Ding, Hsieh and Sharpnack, 2021) and Langevin Monte Carlo methods (Xu et al., 2022) could be adapted for use in zero-inflated models. These methods promise more computationally efficient algorithms, a crucial factor in large-scale and real-time applications. Furthermore, a straightforward extension would be differentiating the feature vectors of the zero component and the non-degenerate distribution component (Poisson or NB) in ZIP and ZINB models. This modification would allow the models to more accurately reflect real-world scenarios where these components are likely to differ.
{acks}
[Acknowledgments] Xueqing Liu is supported by PhD student scholarship from the Duke-NUS Medical School, Singapore. Lauren Bell is supported by a PhD studentship funded by the MRC Network of Hubs for Trials Methodology Research (MR/L004933/2- R18). Bibhas Chakraborty would like to acknowledge support from the grant MOE-T2EP20122-0013 from the Ministry of Education, Singapore.
@@ A.1 Proof of Bayesian regret for TS-NB
$\leq$ $\tau+\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[\Delta_{A_{t}}\right]$
$=$ $\tau+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[R_{t}\left(A_{t}^{*},\beta^{*}\right)-U_{t}\left(A_{t}^{*},\hat{\beta}_{t}\right)\right]}_{\mathcal{R}_{\text{Bayes }}^{(1)}(T)}+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[U_{t}\left(A_{t},\hat{\beta}_{t}\right)-R_{t}\left(A_{t},\beta^{*}\right)\right]}_{\mathcal{R}_{\text{Bayes}}^{(2)}(T)}.$
@@
where $\mathcal{O}(1)$ comes from the failure event of the concentration of $\hat{\beta}_{t}$ in Lemma A.1. According to Lemma 2 in Li, Lu and Zhou (2017), as Assumption 1 in the main manuscript and $\lambda_{\min}(G_{\tau})\geq 1$ are satisfied, we have
@@
$\mathcal{R}_{\text{Bayes}}(T)\leq\tau+\mathcal{O}(1)+\left[2sY_{\max}\kappa_{\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)}.$
For completeness, we set $\tau=\sqrt{d}$ to derive the regret bound presented in Theorem 5.1. Since Algorithm 1 does not incorporate the initial exploration state for parameter estimation, it is feasible to optimize the choice of $\tau$ within the regret bound. ∎
@@ A.2 Proof of Bayesian regret for TS-ZIP
$\leq$ $\tau+\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[\Delta_{A_{t}}\right]$
$=$ $\tau+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[R_{t}\left(A_{t}^{*},\gamma^{*},\beta^{*}\right)-U_{t}\left(A_{t}^{*},\hat{\gamma}_{t},\hat{\beta}_{t}\right)\right]}_{\mathcal{R}_{\text{Bayes }}^{(1)}(T)}+\underbrace{\sum_{t=\tau+1}^{T}\mathbb{E}_{t}\left[U_{t}\left(A_{t},\hat{\gamma}_{t},\hat{\beta}_{t}\right)-R_{t}\left(A_{t},\gamma^{*},\beta^{*}\right)\right]}_{\mathcal{R}_{\text{Bayes}}^{(2)}(T)}.$
@@
$(\|\phi(A_{t},X_{t})\hat{\gamma}-c_{1}^{\gamma}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}}-\phi(A_{t},X_{t})\gamma^{*}\|$
@@
$\leq$ $\sup_{\|o_{1}\|\leq 1,\|o_{2}\|\leq 1}\left\|\nabla f(\bar{o}_{1},\bar{o}_{2})\right\|(2c_{1}^{\gamma}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}}+2c_{1}^{\beta}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}})$
@@
$\mathcal{R}_{\text{Bayes }}(T)$$\leq$ $\tau+\mathcal{O}(1)+\left[2qY_{\max}\kappa_{1,\min}^{-1}\sqrt{d\log(T/d)+4\log T}\right]\sqrt{2dT\log\left(\frac{T}{d}\right)}$
@@
For completeness, we set $\tau=\sqrt{d}$ to derive the regret bound presented in Theorem 5.2. Since Algorithm 2 does not incorporate the initial exploration state for parameter estimation, it is feasible to optimize the choice of $\tau$ within the regret bound. ∎
@@ B.1 Proof of frequentist regret for TS-NB
$\mathbb{E}_{t}\left[\Delta_{A_{t}}\right]\leq s\left(c_{1}+c_{2}\right)\left(1+\frac{2}{p_{3}-p_{2}}\right)\times\mathbb{E}_{t}\left[\left\|\phi(A_{t},X_{t})\right\|_{G_{t}^{-1}}\right]+p_{2}.$
@@
where $\alpha$ and $c_{2}$ are set as in Lemma A.4. For these settings, $p_{2}\leq 1/T$ and $p_{3}\geq 0.15$. Finally, to bound $\sum_{t=\tau+1}^{T}\left\|\phi(A_{t},X_{t})\right\|_{G_{t}^{-1}}$, we use Lemma 2 in Li, Lu and Zhou (2017). As a result, we have
@@ B.2 Proof of frequentist regret for TS-ZIP
$+(\mathbb{P}_{t}(\bar{E}_{2,t}^{\beta})+\mathbb{P}_{t}(\bar{E}_{2,t}^{\gamma})-\mathbb{P}_{t}(\bar{E}_{2,t}^{\beta})\mathbb{P}_{t}(\bar{E}_{2,t}^{\gamma})).$
@@
By setting $\alpha^{\beta}$, $\alpha^{\gamma}$, $c_{2}^{\beta}$, and $c_{2}^{\gamma}$ as in Lemma B.4, we have $p_{2}^{\beta}\leq 1/T$, $p_{2}^{\gamma}\leq 1/T$, $p_{3}^{\beta}\geq 0.15$, and $p_{3}^{\gamma}\geq 0.15$. Finally, to bound $\sum_{\tau+1}^{T}\|\phi(A_{t},X_{t})\|_{G_{t}^{-1}}$, we use Lemma 2 in Li, Lu and Zhou (2017). As a result, we have
@@ Appendix C Technical Lemmas
Using the definition of $S_{t}(\beta^{*})$ and the fact that $\frac{r}{r+\exp(X_{i,A_{i}}^{\top}\beta^{*})}<1$, we can show that $\|S_{t}(\beta^{*})\|_{G_{t}^{-1}}\leq\|Z_{t}\|_{G_{t}^{-1}}$. To bound $\|Z_{t}\|_{G_{t}^{-1}}$, we can use Theorem 1 in Abbasi-Yadkori, Pál and Szepesvári (2011), which states that if the noise is sub-Gaussian with parameter $\sigma$ (in our case, $\sigma=Y_{\max}$), then with probability at least $1-\delta$, we have
@@
By Lemma 11 in Abbasi-Yadkori, Pál and Szepesvári (2011) and Assumption 1 from the main manuscript, we have $\log\det(G_{t})\leq d\log(T/d)$. By the choice of $\tau$, $\det(G_{\tau})^{-1}\leq 1$. It follows that
@@
Appendix D Additional Details for the Drink Less Study
In this section, we provide additional details on the Drink Less study, as well as additional results on different settings.
D.1 Covariates Information
In the algorithms and environment models, we use a subset of the baseline and contextual data that are associated with the outcome. Table 1 shows the features available in the Drink Less MRT dataset.
Table 1: Summary of variables used in the Drink Less study. The features used in $X_{n,t}$ and $S_{n,t}$ are denoted via a $\checkmark$ in the corresponding column, otherwise X.
Feature Description $X_{n,t}$ $S_{n,t}$
Intercept Constant term $\checkmark$ $\checkmark$
Age Age in years $\checkmark$ X
Gender Male (0), Female (1) $\checkmark$ X
AUDIT score Range 0–40 $\checkmark$ X
Days since download Range 0–30 days $\checkmark$ $\checkmark$
Here, $S_{n,t}$ includes variables that interact with the intervention, potentially modifying the effect of sending a notification. We have included “Days since download” based on insights from another study (Liu et al., 2024).
D.2 Additional results
As stated in the manuscript, we conducted simulations to evaluate the algorithms with $p_{\min}=0$ and $p_{\max}=1$. Figure 7 reveals that without probability clipping, the proposed TS-Count algorithms, notably TS-Poisson and TS-NB, outperform others in terms of regret. The comparison shows minimal differences in regret between scenarios with and without probability clipping. This suggests that, with probability clipping, the proposed algorithms maintain consistent regret bounds when compared to a clipped oracle.
(a) OP
(b) ZIOP
Figure 7: Results of the compared algorithms under the Drink Less study with $p_{\min}=0$ and $p_{\max}=1$. The results are calculated from $100$ replications of the experiment. The solid lines indicate the mean values, while the shaded bands represent the standard error bounds across the independent replications.
(a) OP
(b) ZIOP
Figure 8: Quantile plots of the compared algorithms under the Drink Less study with $p_{\min}=0.01$ and $p_{\max}=0.99$. The results are calculated from $1$ replication of the experiment.
Next, we want to study the individual variation in regret across different users. We utilized the same experimental setup as described in Section 7.3 of the manuscript, employing both overdispersed Poisson and zero-inflated overdispersed Poisson models for reward generation.
In Supplement C, we include two interactive plots illustrating user-specific regret for 349 users across a single experiment replication. The plot named ”regret_per_user_op_clipped.html” represents results from an overdispersed Poisson model, while ”regret_per_user_ziop_clipped.html” is derived from a zero-inflated overdispersed Poisson model.
Furthermore, we depict the distribution of user regret across different quantiles in Figure 8. This figure offers a visual analysis of the spread of regret, providing a deeper understanding of its distribution among users.